OpenStudy (anonymous):

Which ones are rational numbers? -12, 0, 96, 8.76, sqrt15, sqrt4, 9 1/3, -23/50, 1.1717717771...

OpenStudy (perl):

i can give similar examples of rational numbers, and based on this you can answer your question. 2 is rational (because 2 = 2/1) sqrt(9) is rational (because sqrt(9)= 3 = 3/1) .11111... is rational (because repeating decimals are rational and .111...=1/9)